OpenStudy (anonymous):

18.01 Single Variable Calculus Problem Set 1 Question 1A-6b

OpenStudy (anonymous):

Hello, could you tell me what you are having trouble with on this problem?

OpenStudy (phi):

if the problem is to find A and ϕ so that \[ \sin x - \cos x = A \sin (x + \phi) \] use this idea \[a \sin x + b \cos x = \sqrt{a^2+b^2}\left( \frac{a}{\sqrt{a^2+b^2}}\sin x +\frac{b}{\sqrt{a^2+b^2}}\cos x\right) \]

OpenStudy (phi):

and this idea 

OpenStudy (phi):

with those definitions \[ a \sin x + b \cos x = \sqrt{a^2+b^2}\left( \cos(\phi)\sin x +\sin(\phi)\cos x\right) \\=\sqrt{a^2+b^2} \sin(x+\phi)\]